A LED company claims that the average life of the LED light bulbs it


manufactures is 1, 500 hours with a standard deviation of 500 hours. If a random


sample of 40 bulbs is chosen, what is the probability that the sample mean will be:


a. greater than 1, 400 hours?


b. less than 1, 400 hours?

1
Expert's answer
2022-03-28T16:02:12-0400

Population mean "\mu=1,500"

Standard deviation "\sigma=500"

Sample size "n=40"


"Z=\dfrac{\bar x-\mu}{\sigma/\sqrt n}=\dfrac{1400-1500}{500/\sqrt{40}}"


"Z=-1.26"


(a) "\bar x>1400"

"=1\u2212P(z<\u22121.26)"

"=1\u22120.10565"

"=0.89435"


(b) "\bar x<1400"

"=P(z<\u22121.26)"

"=0.10565"